a certain grade egg must weigh at least 2 oz. If the weights of eggs are normally distributed with a mean of 1.5oz and a standard deviation of 0.4oz, appromately how many eggs in a random sample of 29 dozen would you expect to weigh more than 2 oz?
Please show me what formula to use!

Answer by stanbon(75887)

a certain grade egg must weigh at least 2 oz. If the weights of eggs are normally distributed with a mean of 1.5oz and a standard deviation of 0.4oz, appromately how many eggs in a random sample of 29 dozen would you expect to weigh more than 2 oz?
------------
z(2) = (2-1.5)/0.4 = 0.5/0.4 = 1.25
P(x > 2) = P(z > 1.25) = 0.1056
---
# more than 2 oz = 0.1056*29*12 = 36.7 eggs
------------------
Cheers,
Stan H.